# Generative AI models enable efficient and physically consistent sea-ice simulations

### v1

@@ Abstract
Sea ice is governed by highly complex, scale-invariant, and anisotropic processes that are challenging to represent in Earth system models. While advanced numerical models have improved our understanding of the sea-ice dynamics, their computational costs often limit their application in ensemble forecasting and climate simulations. Here, we introduce GenSIM, the first generative AI-based pan-Arctic model that predicts the evolution of all relevant key properties, including concentration, thickness, and drift, in a 12-hour window with improved accuracy over deterministic predictions and high computational efficiency, while remaining physically consistent. Trained on a long simulation from a state-of-the-art sea-ice–ocean system, GenSIM robustly reproduces statistics as observed in numerical models and observations, exhibiting brittle-like short-term dynamics while also depicting the long-term sea-ice decline. Driven solely by atmospheric forcings, we attribute GenSIM’s emergent extrapolation capabilities to patterns that reflect the long-term impact of the ocean: it seemingly has learned an internal ocean emulator. This ability to infer slowly evolving climate-relevant dynamics from short-term predictions underlines the large potential of generative models to generalise for unseen climates and to encode hidden physics.
@@
Sea ice regularises the Earth system by driving thermohaline circulation and modulating albedo. At the mechanical scale, its complex dynamics arise from the transition of a fluid flow to a fractured, discrete continuum, marked by scale invariance, multi fractality, and anisotropy [1, 2]. These processes are captured in numerical models through rheologies which govern the simulated mechanical response to external forcings coming from the atmosphere and ocean such as winds, currents, and waves [3]. In particular, a new branch of sea-ice models with brittle-class rheologies [4, 5, 6, 7, 8] recently introduced a damage variable to parametrise the subgrid fracturing of sea ice and to reproduce the spatial and temporal properties of its dynamics across scales as observed from high-resolution satellite observations [9, 6, 10]. Yet, these rheologies, typically implemented with advanced numerical techniques such as Lagrangian finite elements [5], increase computational cost [11, 7], thereby complicating ensemble forecasts and long-term climate projections.
@@
This highlights a long-standing dilemma: sea-ice models that faithfully reproduce the observed dynamics require either advanced rheologies or kilometre-scale resolutions [10], whereas computationally efficient models sacrifice realism. Computational gains are offered by surrogate models that leverage deep neural network, however, they remain limited to predictions of the sea-ice thickness [12] or have narrow applications [13, 14, 15]. Crucially, these models suffer from a loss of small-scale information [12, 16], smoothing [17], and reduced physical consistency [18] compared to physics-based numerical models. Moreover, their auto-regressive predictions risk an error accumulation [19] leading to instabilities in long-term simulations [20].
@@
Generative diffusion and flow matching models [21, 22, 23, 24, 25] remedy these limitations by learning to produce diverse realisations conditioned on the current states and external forcings. This enables applications from ensemble generation [26, 27], over weather prediction [28] and downscaling [29], to climate applications [30, 31, 32]. When applied to sea-ice predictions, these models can refine sea-ice concentration forecasts [33], generate sea-ice states [34], and predict the sea ice on regional scales [35]. Nonetheless, their iterative generation remains unconstrained by physical bounds, and their inherent computational demands hamper adoption to global-scale predictions.
@@
We resolve this dilemma with GenSIM, the first generative AI-based pan-Arctic sea-ice model that predicts all relevant sea-ice states across spatial scales of $10\,\text{km}$$100\,\text{km}$ and temporal scales from $12\,\text{hours}$ to multi-decadal climate trends. GenSIM simultaneously achieves: (i) a high computational efficiency–being up to $50\times$ faster than the reference sea ice-ocean configuration neXtSIM‑OPA [36, 5, 37]–and (ii) an unprecedented level of physical consistency for AI-based sea-ice models, as exemplified through the strict enforcement of non-negativity for the sea-ice thickness. Instantiated as auto-regressive flow matching model, GenSIM employs a scale-aware transformer architecture and exploits physical localisation by domain decomposition for efficient predictions. Trained on 20 years of neXtSIM-OPA output [36] and validated against numerical models and satellite observations, GenSIM reproduces realistic the large-scale sea-ice evolution over 30 years, including the warming-climate-driven sea-ice decline, the recent hiatus due to internal variability [38], and multifractal brittle-like short-term dynamics. Despite being forced only by the atmosphere, GenSIM implicitly learns ocean-like patterns in its dynamics and thermodynamics–this emergent capability overcomes the conventional requirement for explicit ocean data. Thus, a generative AI model trained for short-term forecasting can robustly perform long-term multi-decadal predictions and implicitly capture multi‑scale physical processes, thereby removing a bottleneck in AI-based Earth system modelling.
1 Sea-ice modelling with GenSIM
Figure 1: Schematic of the efficient Arctic-wide generative AI-based sea-ice model GenSIM. Based on flow matching, GenSIM jointly predicts all important sea-ice states for a lead time of 12 hours. (a) Conditioned on initial states $\mathbf{x}_{t}$ and external atmospheric forcings $\mathbf{F}_{t:t+12\,\text{h}}$, the model iteratively maps fields of pure noise $\mathbf{z}^{0}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$ to the predicted residual $\mathbf{z}^{1}$, which, added to the initial conditions, results into the prediction $\mathbf{x}_{t+12\,\text{h}}$ 12 hours later. (b) Within an iterative flow step, here at pseudo time $\tau$, the full Arctic domain is decomposed into overlapping subdomains. A geophysical transformer processes all subdomains in parallel and predicts the velocity of the flow, which is then used to update the state for the next pseudo time $\tau^{\prime}$. The embedding for modulating the transformer includes information about the pseudo time, data augmentation labels, and used resolution.
@@ 1 Sea-ice modelling with GenSIM
GenSIM is a sea-ice model developed to generate 12-hour pan-Arctic forecasts of sea-ice thickness, concentration, damage, velocity components, and snow-on-ice thickness. In addition to the initial conditions, the model accounts for external forcings from the atmosphere: the 2-metre temperature and specific humidity, and 10-metre zonal and meridional wind velocities. To extend the forecast window beyond the trained 12 hours, GenSIM is implemented as an autoregressive system, where each prediction serves as initial condition for the subsequent time step, similar to prediction models for the atmosphere [28]. Natively operating on a curvilinear $1/4^{\circ}$ mesh (approximately 12 km resolution), GenSIM makes forecasts in a few seconds and needs below $4\,\text{GB}$ memory, which enables forecast ensembles and decadal-long simulations on a single consumer-grade GPU.
@@
Instantiated as flow matching model [24] and conditioned on the initial conditions, $\mathbf{x}_{t}$, and forcings, $\mathbf{F}_{t:t+12\,\text{h}}$, GenSIM produces forecasts by iteratively mapping samples from a Gaussian distribution, $\mathbf{z}^{0}$, to tendency fields, $\mathbf{z}^{1}$, as schematically depicted in Fig. 1a. Adding these tendencies to the initial conditions yields the 12-hour forecast, $\widehat{\mathbf{x}}_{t+12\,\text{h}}$. Crucially, varying the initial noise sample allows GenSIM to produce diverse forecasts from identical initial conditions and forcings.
Integrated in pseudo time from $\tau=0$ (pure noise) to $\tau=1$ (clean tendency), the flow and the underlying dynamical system are defined by its velocity, $\widehat{\mathbf{v}}$. A trained neural network approximates the velocity to propagate the state to the next pseudo-time step, $\mathbf{z}^{\tau^{\prime}}$, given the intermediate states $\mathbf{z}^{\tau}$, the initial conditions and forcings, and any relevant embeddings.
Observing the parallels between this generative flow and traditional numerical methods [27], and leveraging the localised physical sensitivity of the tendencies, we introduce a domain decomposition strategy to enhance the computational efficiency at global scale [39, 32]. Instead of processing global fields, the neural network predicts velocities within decomposed subdomains (Fig. 1b). To facilitate the communication between subdomains when integrating the flow, the inputted subdomains overlap, while the velocity prediction itself remains local. This decomposition allows for fully parallel processing and enables linear scaling with the number of global grid points, i.e., the number of subdomains is doubled when the number of global grid points is doubled.
The neural network adopts a scale-aware, mesh-independent transformer architecture, designed to explicitly encode spatial interactions and to accommodate masked land cells. Initially, the network converts the input into tokens through a linear $2\times 2$ patching [40]. These tokens are processed by a series of transformer blocks: self-attention blocks capture spatial dependencies with a learnable localisation that depends on relative distances [41, 42], while light-weighted multi-layer perceptrons (MLPs) non-linearly combine the resulting spatial representations. Derived from the pseudo time and auxiliary labels such as the resolution, a context embedding modulates the input to the attention and MLP layers. After $8$ transformer blocks, the tokens are mapped to local velocity predictions via a detokenisation layer.
@@
Sea ice is physically bounded (e.g., the non-negativity of the thickness), and we constrain the generative flow to follow the appropriate bounds [43]. To expose the neural network to bounds during training, we introduce censored flow matching based on variational flow matching [44] and censored Gaussian distributions [34]. The resulting loss function adjusts the usual mean-squared error when encountering bounded data, e.g., over open ocean. There, the loss drives the model towards more extreme values to increase the probability that the bound is exceeded, as concisely exposed in Section 4.3 and detailed derived in Appendix B. This way, the model effectively accounts for the bounds while predicting sea ice.
@@
We trained the model on 20 years (1995–2014) from a long simulation with the coupled sea-ice–ocean neXtSIM-OPA system [36] and forced by the atmosphere with the ERA5 reanalysis [45]. To leverage the scale invariance of sea ice and to enhance mesh and resolution independence, we augmented the data by averaging over randomly drawn kernel sizes up to $8\times 8$. After averaging, the domain decomposition is implemented during training by randomly slicing subdomains from the pan-Arctic domain to match the subdomain size used during sampling ($80\times 80$ grid points, including $8$ points overlap in each direction). This efficiently reduced the training cost and enabled us to train the model on fewer GPUs [46]. Finally, we employed distribution augmentation with random flipping and rotations [47, 35] to promote the model’s ability to generalise.
Figure 2: Overview of GenSIM’s short-term forecast qualities. Shown are: (a) the improvement in the root-mean-squared error (RMSE) of GenSIM’s ensemble mean compared to the deterministic model and the ratio between the ensemble standard deviation to the RMSE of GenSIM’s ensemble mean, both are averaged over all ice-covered grid points (sit=thickness, sic=concentration, sid=damage, siu=$u$-velocity, siv=$v$-velocity, snt=snow-on-ice thickness); (b) the 10-day RMSE improvement of GenSIM’s ensemble mean to the deterministic model split up into regions, averaged across all six predictive variables after normalizing with their climatology; (c) the 10-day spectrum of the sea-ice thickness in the central Arctic, averaged in log-space across all forecasts; (d) accuracy of the 10-day forecast for the sea-ice extent with the median accuracy for the ensemble members. The used regions for (b) and (c) are depicted in Fig. 6.
We test GenSIM’s short-term forecasting qualities in a perfect model experiment for 2016–2018, where the simulation of neXtSIM-OPA is our truth, against which we evaluate. Initialised from the simulation and driven by ERA5 forcings, we start a 10-day forecast once a week, effectively performing $153$ initialisations. Running an ensemble of $16$ members from the same initial conditions and forcings, we compare the performance against persistence forecasts and against a deterministic deep learning model, based on a U-ViT architecture [34] and trained with a mean-squared error loss for the same 12-hour lead time.
Using the root mean-squared error (RMSE) as metric, GenSIM’s ensemble mean outperforms these baselines (Fig. 2a). Across all six predicted variables and all regions, GenSIM improves the scores, reaching an improvement of around $20\,\%$ compared to the deterministic forecasts. In the central Arctic, which is mostly year-long covered by sea ice, GenSIM achieves substantial improvements, as well as in regions with marginal ice zones (MIZ, Fig. 2b), though small deficiencies in the sea-ice concentration during the change from melting to refreezing remains (see also Fig. 7 and Fig. 8). These results are striking: while the deterministic model is specifically trained to minimise the RMSE, GenSIM–designed as generative flow–achieves an indirect reduction. Derived from independent predictions, GenSIM’s ensemble mean is a true statistical average, explaining why it can outperform the deterministic model [18]. Although well-tuned after 12 hours, GenSIM’s ensemble becomes seemingly underdispersive with increasing lead time, leading to typical spread-skill ratios of $0.8$ after 10 days. This underdispersion could be remedied by injecting more stochasticity [28] or tuning on the validation dataset [48] from which we refrain.
Like the geophysical model neXtSIM, GenSIM retains the energy at all scales even after a 10-day forecast (Fig. 2c), while the ensemble mean forecast and the deterministic model lose small-scale energy. This loss of small-scale information effectively smooths the forecast out [12, 35], leading to physical inconsistencies [18] and potentially incurring instabilities in long-term simulations. GenSIM additionally predicts the sea-ice extent more accurately, which becomes especially evident during summer and autumn (Fig. 2d). GenSIM seems to better represent the sea-ice edge and marginal ice zone (MIZ), which we partially attribute to the generative model and partially to the novel censored loss function that explicitly incorporates the physical bounding of variables. These results confirm previous findings for a regional generative sea-ice model [35] and validate that the model works on short time scales. In conclusion, GenSIM outperforms the deterministic model and resolves the problems with the loss of small-scale information, like similarly found for the atmosphere [28], while better constraining the MIZ.
2 Realistic long-term simulations through generative short-term predictions
@@
To analyse the long-term behaviour of GenSIM, we perform a long simulation, similar to the neXtSIM-OPA simulation that was used to train our model. Starting with the same initial conditions as neXtSIM-OPA and forced by ERA5, GenSIM is run autoregressively over 1995–2024 with the first five years used as spin-up. Since the model is trained for short-term predictions on data up to 2015, we test GenSIM’s stability over 2000–2014, while we focus on statistical averages and its extrapolation capabilities on the subsequent period. GenSIM’s simulation is benchmarked against the original neXtSIM-OPA simulation and the PIOMAS reanalysis product [49]. For the estimation of the volume, we additionally use a satellite product derived from CryoSat2, and Soil Moisture and Ocean Salinity (SMOS), called CS2/SMOS [50].
Figure 3: Overview over GenSIM’s long-term projection qualities. (a) The simulated monthly averaged sea-ice volume (blue line) compared to the simulated volume of neXtSIM (red dotted line) and the retrieved volumes from PIOMAS (grey dashed line) and the merged CryoSAT2-SMOS product (CS2/SMOS, black crosses). (b, c) The change in the winter thickness between 2020-2024 to 2000–2004 for GenSIM and PIOMAS. (d, e) The winter sea-ice thickness averaged between 2015–2018 for GenSIM and neXtSIM. (f–i) Snapshots of GenSIM’s simulation for 2024-12-25 03:$0$, shown are the sea-ice concentration (f), instantaneous shear rate (g), instantaneous divergence rate (h), and the change in the sea-ice thickness to 2024-12-25 15:$0$ (i).
GenSIM with its autoregressive steps is remarkably stable over 30 years, even in a warming climate (Fig. 3a): the simulation successfully reproduces observed seasonal cycles and closely aligns with the results from neXtSIM and PIOMAS. Yet, GenSIM is exposed to a small positive bias, especially during summer, attenuating the Arctic sea-ice regime shift post-2007 [51]. As a consequence, GenSIM underestimates the climate change impact for the period 2000–2018 with a linear trend of $224\,\text{km}^{3}\,\text{year}^{-1}$ compared to $252\,\text{km}^{3}\,\text{year}^{-1}$ for neXtSIM and $370\,\text{km}^{3}\,\text{year}^{-1}$ for PIOMAS. Nevertheless, as caused by internal variability, GenSIM successfully exhibits the recent slowing of sea-ice decline consistent with satellite observations and reanalysis datasets [38].
By comparing the spatial distribution of change in average sea-ice thickness between 2000–2004 and 2020–2024 for GenSIM and PIOMAS, the underestimation of the trend becomes especially evident in the central Arctic (Fig. 3b and c). Nevertheless, GenSIM’s sea-ice thickness distribution closely follows neXtSIM’s for 2015–2018 (Fig. 3d and e), even though a small bias towards thicker sea ice persists. The reduced, and in some areas reversed, climate change impact on sea ice surrounding Greenland, potentially attributable to the influx of cooler melt water, is captured by GenSIM. This suggests that GenSIM has learned the well-established relationship between atmospheric near-surface temperature and sea ice from the 2-metre temperature used as external forcing.
The fields generated by GenSIM remain physically consistent, even after $30$ years of simulation (Fig. 3f–i). We can track the effect of the short-term dynamics throughout all fields: a reduction in the sea-ice concentration suggests a weakening of sea ice, making it more vulnerable against stress induced by atmospheric winds. The external stress results into a shear and divergent deformation, which are as a diagnostic extracted from the predicted sea-ice velocities. This deformation leads to a redistribution of sea ice with linking divergence to a thinning and convergence to a thickening of sea ice. Although only trained for 12-hour predictions, GenSIM’s largely consistent trend with numerical models and satellite observations, despite a small positive bias, and its physical consistency demonstrates its potential for projecting towards future climate scenarios.
Figure 4: Fractal scaling of GenSIM’s dynamics. (a–c) Triangulated total deformation rates as inferred from RGPS observations (a), from advecting with GenSIM’s predicted velocities (b), and from advecting with neXtSIM’s predicted velocities (c) for the period 2007-02-01 03:00–2007-02-07 15:$0$. The deformation rates are estimated at the original data resolution with a nominal time scale of 3 days. (d) Distribution of instantaneous total deformation rates for GenSIM and neXtSIM at the nominal resolution of around $12.5\,\text{km}$. (e) First three distributional moments of the instantaneous total deformation rates with a decreasing coarse-grained resolution. (f) Scaling coefficients of the multifractal structure function $\beta(q)$ from the relation $\big{\langle}\dot{\epsilon}^{q}_{\mathrm{tot}}(L)\big{\rangle}\sim L^{-\beta(q)}$ with $L\in[12.5\,\text{km},400\,\text{km}]$ for the first three distributional moments $q$ of the instantaneous total deformation rates. The uncertainty in the structure function shows the minimum and maximum in the coefficients estimated based on differences between two successive spatial scales [9, 35]. In (d)–(f), the total deformation rates are estimated based on rectangles defined through the curvilinear mesh in the central Arctic (see also Fig. 6) for January, February, March during the period 2015-01-01 03:00–2018-03-31 15:$0$.
The brittle Bingham-Maxwell rheology [6] employed in neXtSIM is designed to parametrise the short-term subgrid dynamics of sea ice as observed by satellites. The multifractal scaling of GenSIM’s short-term dynamics is analysed based on the sea-ice deformation for the winter in the period 2015–2018, where no high-resolution observations of the sea-ice motion are available.
GenSIM demonstrates its strong skill to reproduce brittle key characteristics of observed Arctic sea-ice deformation from RGPS, while performing comparably to neXtSIM. Although exhibiting more diffuse structures due to its lower resolution (the median resolution of RGPS is around $10\,\text{km}$), GenSIM successfully captures the spatial organisation and intensity of the deformation features from RGPS (Fig. 4a–c) and neXtSIM, including linear kinematic features and high-strain bands.
The scale invariance of sea ice is exposed by the spatial scaling of the distributional moments from the deformation rates with the relation $\left\langle\dot{\epsilon}^{q}_{\mathrm{tot}}(L)\right\rangle\sim L^{-\beta(q)}$ [9, 6]. Here, GenSIM maintains the behaviour of the sea-ice dynamics as modelled by neXtSIM: the heavy-tailed distribution of total deformation rates from GenSIM (Fig. 4d) mirrors neXtSIM’s with a minor underestimation at the highest rates. Moreover, the spatial scaling of GenSIM’s dynamics closely aligns with neXtSIM’s across orders of magnitude (Fig. 4e). The scaling coefficient for the mean ($\beta(q=1)=0.13$) is in the range found in other studies [9]. Inferred from the change in the total deformation rate, the multifractal structure function, as change of the scaling coefficient with the distributional moment, shows that GenSIM exhibits a similar multifractal scaling to neXtSIM (Fig. 4f).
@@
Trained to predict the sea ice over 12 hours with external atmospheric forcings, GenSIM constrains the short-term dynamics well, producing realistic scale-invariant and anisotropic brittle-like features. However, long-term dynamics and especially the thermodynamics are difficult to learn due to their diminishing signal-to-noise ratio and the missing ocean forcings.
Figure 5: GenSIM exhibits long-term ocean-like patterns. (a–c) Volume change within 12 hours for (a) GenSIM, (b) neXtSIM, and (c) net bottom growth - bottom melt from neXtSIM, averaged for 2015–2018. The dynamical component caused by advection has been removed from the changes. (d–f) Speed (background) and streamlines (black lines) for (d) sea ice in GenSIM, (e) sea ice in neXtSIM, and (f) ocean surface from the neXtSIM-OPA simulation averaged for 2015–2018; the speed for the dark blue circled position is shown in (g). (g) Time series of speed in the Fram strait with a year-long averaging filter for long-term tendencies, the shaded area indicates the averaging period for (a–f); note that the atmospheric speed has been scaled down to $3\,\%$ of its original speed for visualisation purposes.

### v2

# Generative AI models capture realistic sea-ice evolution from days to decades

@@ Abstract
Sea ice plays an important role in stabilising the Earth system. Yet, representing its dynamics remains a major challenge for models, as the underlying processes are scale-invariant and highly anisotropic. This poses a dilemma: physics-based models that faithfully reproduce the observed dynamics are computationally costly, while efficient AI models sacrifice realism. Here, to resolve this dilemma, we introduce GenSIM, the first generative AI model to predict the evolution of the full Arctic sea-ice state at 12-hour increments. Trained for sub-daily forecasting on 20 years of sea-ice–ocean simulation data, GenSIM makes realistic predictions for 30 years, while reproducing the dynamical properties of sea ice with its leads and ridges and capturing long-term trends in the sea-ice volume. Notably, although solely driven by atmospheric reanalysis, GenSIM implicitly learns hidden signatures of multi-year ice–ocean interaction. Therefore, generative AI can extrapolate from sub-daily forecasts to decadal simulations, while retaining physical consistency.
@@
Sea ice stabilises the Earth system by regulating thermohaline circulation and modulating albedo. Its mechanical dynamics arise from the transition of a fluid flow to a fractured, discrete continuum, marked by scale invariance, multi fractality, and anisotropy (?, ?). These processes are captured in physics-based numerical models through rheologies that describe how the sea ice deforms under external forcings such as winds, currents, and waves (?). Recent sea-ice models with brittle rheologies (?, ?, ?) use a damage variable to parametrise subgrid fracturing and to reproduce the multi-scale dynamics observed in high-resolution satellite data (?, ?, ?). Although this increases their realism, these models typically require advanced numerical techniques such as Lagrangian finite elements (?), which raise their computational costs and complicate ensemble forecasts and long-term simulations (?, ?).
@@
This highlights a long-standing dilemma: sea-ice models that faithfully reproduce the observed dynamics require either kilometre-scale resolutions or brittle rheologies (?), whereas computationally efficient models sacrifice realism. In recent years, computational gains are offered by surrogate models that leverage deep neural network. For sea ice, most approaches so far have focused on univariate predictions of sea-ice thickness (?) or sea-ice concentration and extent (?, ?, ?), missing the full picture of the sea-ice state. Crucially, trained as deterministic models, they suffer from a loss of small-scale information and smoothing (?, ?), and reduced physical consistency (?) compared to physics-based models. Moreover, their auto-regressive predictions risk an error accumulation (?) leading to instabilities in long-term simulations (?).
@@
Generative diffusion and flow matching models (?, ?, ?, ?, ?) remedy these limitations by learning to produce diverse state realisations. Conditioned on the current state of the system and external forcings, this enables applications from ensemble generation (?, ?, ?), over weather prediction (?) and downscaling (?), to climate emulation (?, ?). When applied to sea-ice predictions, these models can refine sea-ice concentration forecasts (?), generate sea-ice states (?), and predict sea ice on regional scales (?). Nonetheless, their computational demands hamper adoption to global-scale predictions, and their iterative generation remains unconstrained by physical bounds such as non-negativity of the sea-ice thickness.
@@
We resolve this dilemma with GenSIM, the first generative AI-based sea-ice model that predicts the full sea-ice state across spatial scales from $10\,\text{km}$ to the Arctic basin and temporal scales from $12\,\text{hours}$ to multi-decadal climate trends. Instantiated as an auto-regressive flow matching model, GenSIM utilises a scale-aware transformer architecture and exploits physical localisation by domain decomposition for efficient predictions, being up to 50 times faster than the reference physics-based neXtSIM-OPA configuration (?). Trained on 20 years of neXtSIM-OPA output, GenSIM’s novel censored loss function enforces physical bounds during training. These advances allow GenSIM to make realistic predictions over 30 years, reproducing multifractal brittle-like short-term dynamics, the decades-long sea-ice decline, and the recent hiatus due to internal variability (?), as validated against physics-based models and satellite observations Although forced solely by the atmosphere, GenSIM learns to extrapolate by implicitly encoding the multi-year interactions between sea ice and ocean. Trained for sub-daily forecasting, GenSIM can robustly perform multi-decadal predictions and capture multi-scale physical processes, thereby effectively removing the common inability of simultaneously learn multiple scales in AI-based Earth system models.
Sea-ice modelling with GenSIM
@@
GenSIM is a sea-ice model developed to generate 12-hour forecasts of sea-ice thickness, concentration, damage, drift components, and snow-on-ice thickness over the full Arctic basin. The model is driven by its initial conditions and external forcings from the atmosphere: the 2-metre temperature and specific humidity, and 10-metre zonal and meridional wind. To extend the forecast window beyond the trained 12 hours, GenSIM is implemented as an autoregressive system, where each prediction serves as initial condition for the subsequent 12 hours, similar to prediction models for the atmosphere (?). Natively operating on a curvilinear $0.25^{\circ}$ mesh (approximately 12 km resolution), GenSIM makes 12 hour forecasts in a few seconds and needs below $4\,\text{GB}$ memory, which enables large ensembles and decadal-long simulations on a single consumer-grade GPU (here, an Nvidia RTX A6000 Ampere with $48\,\text{GB}$ VRAM).
@@
Instantiated as flow matching model (?) and conditioned on the initial conditions, $\mathbf{x}_{t}$, and forcings, $\mathbf{F}_{t:t+12\,\text{h}}$, GenSIM produces forecasts by iteratively mapping samples from a Gaussian distribution, $\mathbf{z}_{0}$, to tendency fields, $\mathbf{z}_{1}$, as schematically depicted in Fig. 1a. Added to the initial conditions, these tendencies yield the 12-hour forecast, $\widehat{\mathbf{x}}_{t+12\,\text{h}}$. Crucially, varying the initial noise sample allows GenSIM to produce an ensemble of forecasts from identical initial conditions and forcings, representing model uncertainty.
@@
Sea ice is physically bounded (e.g., non-negativity of the thickness). To follow these bounds, we introduce censored flow matching based on variational flow matching (?) and censored Gaussian distributions (?), effectively constraining the generative flow. The resulting loss function adjusts the usual mean-squared error when encountering bounded data. There, the loss drives the model towards more extreme values to increase the probability that the bound is exceeded as exposed in the mathematical framework (Supplementary Text).
Integrated in pseudo time from $s=0$ (pure noise) to $s=1$ (clean tendency), the flow and the underlying dynamical system are defined by its generative velocity, $\widehat{\mathbf{v}}_{s}$. A trained neural network approximates the generative velocity to propagate the state to the next pseudo-time step, $\mathbf{z}_{s^{\prime}}$, given the intermediate states $\mathbf{z}_{s}$, the initial conditions, forcings, and any relevant embeddings as derived from the pseudo time, augmentation labels, and the used resolution.
Considering the similarities between this generative flow and traditional numerical methods (?), and leveraging the localised physical sensitivity of the tendencies, we use a domain decomposition strategy (?, ?) to enhance computational efficiency. Instead of processing global fields, the neural network predicts velocities within decomposed subdomains (Fig. 1b). To foster communication between subdomains when integrating the flow, the inputted subdomains overlap, while the predicted velocity itself remains local. This decomposition enables for fully parallel processing and a linear scaling with the number of global grid points, i.e., the number of subdomains is doubled when the number of global grid points is doubled.
The neural network adopts a scale-aware, mesh-independent transformer architecture, designed to explicitly encode spatial interactions and to accommodate masked land cells. Initially, through a linear tokenizer with a $2\times 2$ patching (?), the network converts the input into tokens. These tokens are processed by a series of transformer blocks: self-attention blocks capture spatial dependencies with a learnable localisation that depends on relative distances (?, ?), while multi-layer perceptrons (MLPs) non-linearly combine the resulting spatial representations. The embedding modulates thereby the inputs to the attention and MLP blocks. After $8$ transformer blocks, the tokens are mapped to local velocities via a detokenisation layer.
@@
We trained the model on 20 years (1995–2014) from a long simulation with the coupled sea-ice–ocean neXtSIM-OPA system (?) and forced by the atmosphere with the ERA5 reanalysis (?). To leverage the scale invariance of sea ice and to enhance mesh and resolution independence, we augmented the data by averaging over randomly drawn kernel sizes up to $8\times 8$. The domain decomposition is implemented during training by randomly slicing subdomains from the pan-Arctic domain to match the sampling subdomain size ($80\times 80$ grid points, including $8$ points overlap in each direction). This efficiently enabled us to train the model on eight GPUs only and cut the training cost to around two days (?). Finally, we employed distribution augmentation (?) with random flipping and rotations to promote the model’s ability to generalise.
Before applying GenSIM to decadal long-term simulations, we evaluated its performance in emulating neXtSIM-OPA on held-out data from 2016–2018 (Supplementary Text, figure S2, figure S3, and figure S4). GenSIM outperforms a deterministic baseline for this period, reducing in its mean the root-mean-squared error by around $20\,\%$, matching the energy spectrum from neXtSIM-OPA for individual members, and improving the representation of the marginal ice zone. Additionally, GenSIM’s ensemble exhibits typical spread-skill ratios around $0.8$, remaining just slightly underdispersive. These results make GenSIM a major step towards realistically emulating physics-based sea-ice models and improving upon deterministic models.
Realistic decade-long simulations
@@
To analyse the long-term behaviour of GenSIM, we perform a 30 years long simulation with a single ensemble member, similar to the neXtSIM-OPA simulation that was used to train our model. Starting from the same initial conditions as neXtSIM-OPA and forced by ERA5, GenSIM is run autoregressively over the period 1995–2024 with the first five years used as spin-up and the last ten years as extrapolation period.
Driven by the mechanical response to external atmospheric forcings, sea ice exhibits a cycle of processes from opening of leads, over healing, to ridging. Even though GenSIM has no processes explicitly encoded, it successfully captures their impact in space and time (Fig. 2): (I) a small reduction in the sea-ice concentration suggests a weakening of sea ice, making it more vulnerable against stress induced by atmospheric winds. The external stress results into a divergent deformation, which leads to ice thinning. This thinning amplifies the response such that larger leads are opened. (II) After the external forcing is removed or when the stress is sufficiently decreased, opened leads are healed, resulting in a subsequent increase in concentration. (III) When linked to dynamical ridging, this healing is often accompanied by convergence, where sea ice is piled up, resulting into a strong thickening of ice. The fact that we can track the impact of these processes in the predictions suggests that GenSIM remains physically consistent even after nearly 30 years of simulation.
This physical consistency demonstrates the strengths of generative flow models in learning high-dimensional correlations and mechanical processes in space and across variables. Although the tendency fields are independently generated from Gaussian noise for each prediction time, they are linked in time through auto-regressive conditioning. Moreover, as the tendencies are added up over time, GenSIM yields temporal consistency as required to exhibit such realistic cycles of damaging and healing. Consequently, the auto-regressive application of generative flows by predicting tendencies is likely key to achieve this physical consistency in space and time.
@@
Trained to predict the sea ice over 12 hours with external atmospheric forcings, GenSIM constrains the short-term dynamics well, producing realistic scale-invariant and anisotropic brittle-like features (Supplementary Text,  figure S5). Long-term dynamics and especially the thermodynamics are however difficult to learn due to their diminishing signal-to-noise ratio on sub-daily scales and, importantly, the missing ocean forcings.
To benchmark the thermodynamical properties, we compare GenSIM’s simulated sea-ice volume to the original neXtSIM-OPA simulation and the PIOMAS reanalysis product (?). As observational references, we additionally use satellite products derived from CryoSat2, and Soil Moisture and Ocean Salinity (SMOS), called CS2/SMOS (?), as well as IceSAT2 observations (?).
Remarkably, GenSIM with its autoregressive steps is stable over 30 years, even in a warming climate (Fig.  3a): the simulation successfully reproduces the observed seasonal cycle and closely aligns with the results from neXtSIM and PIOMAS (Fig. 3b). GenSIM shows a small positive bias, especially during summer (September), attenuating the Arctic sea-ice regime shift post-2007 (?). This bias is primarily due to an overestimation of the sea‑ice area during summer (Supplementary Text, figure S7). As a consequence, GenSIM underestimates the climate change impact for the period 2000–2018 with a linear trend of $187\pm 10\,\text{km}^{3}\,\text{year}^{-1}$ compared to $215\pm 9\,\text{km}^{3}\,\text{year}^{-1}$ for neXtSIM and $329\pm 12\,\text{km}^{3}\,\text{year}^{-1}$ for PIOMAS (only uncertainties from the regresssion itself are included). Overestimating the observed minimum in 2012, GenSIM and neXtSIM reach their lowest September estimates in 2017 and 2018. Nonetheless, GenSIM successfully predicts the more recent slowing of sea-ice decline (trend estimate: $104\pm 29\,\text{km}^{3}\,\text{year}^{-1}$ for the period 2015–2024) due to Earth system’s internal variability, consistent with satellite observations and reanalysis datasets (?) (Supplementary Text, figure S6).
By comparing the spatial distribution in averaged winter sea-ice thickness for the periods 2000–2004 and 2020–2024 for GenSIM and PIOMAS, the underestimation of the trend becomes especially evident in the central Arctic (Fig. 3c and d). Yet, GenSIM’s spatial thickness distribution closely follows neXtSIM’s for 2015–2018 (Fig. 3e and f), despite a persisting small bias towards thicker sea ice. Even though GenSIM is exposed to biases in its volume estimates and trends, these results suggest that GenSIM has learned the seasonal character of sea ice and the long-term impact of climate change, as also seen in the diagnosed sea-ice age (Supplementary Text, figure S8).
The neXtSIM-OPA system (?) explicitly accounts for sea-ice–ocean interactions by coupling the sea-ice model neXtSIM (?, ?) to the ocean component of the NEMO modelling framework (?). Being solely forced by the atmosphere, GenSIM contrastingly neglects an explicit representation of these interactions. Nevertheless, GenSIM captures the long-term decline in sea-ice volume as driven by thermodynamics, with only a small bias compared to the neXtSIM-OPA simulation. Consequently, we need to understand how GenSIM reproduces these long-term impacts so well.